# A composition of simplified physics-based model with neural operator for trajectory-level seismic response predictions of structural systems

### v1

@@ 5.1.3 Parametric study
A parametric study is conducted to evaluate neural operator model sensitivity to training dataset size and the number of Fourier kernel layers. Figures 7 and 8 show relative $\mathcal{L}^{2}$ errors across 10 runs for varying configurations. These experiments are conducted for C-PhysFNO using ELS as $\mathcal{H}_{p}$, without postprocessing. While the performance of the standard FNO improves with increased data availability and deeper architectures, the C-PhysFNO consistently achieves lower error across all cases. Notably, the performance gap is most pronounced under data-limited conditions, highlighting the advantage of integrating domain knowledge through physics-based intermediate representations.
@@ 5.2.1 Structural model and ground motion dataset
This example investigates the seismic response of the Auburn Ravine bridge, a prestressed concrete box-girder structure comprising six spans and five bents, each supported by two circular piers [45]. A detailed finite element model is developed in OpenSees [46], incorporating geometric nonlinearity and material nonlinearities in the pier columns and connection regions. The total bridge length is 166.4 m. The piers are modeled using force-based beam-column elements with fiber-section discretization. Confined and unconfined concrete fibers and reinforcing steel bars represent axial–bending interaction, while shear and torsion are modeled using uniaxial components. The deck is rigidly connected to the piers, which are fixed at their bases. Translational springs are used to model abutment stiffness. The fundamental period of the bridge is $T=0.38$ s. The quantity of interest is the displacement time history at the top of each of the ten piers, denoted by $\boldsymbol{u}(t)=\{u_{1}(t),\dots,u_{10}(t)\}$.
The ground motion dataset consists of spectrum-compatible synthetic accelerograms generated using the frequency-domain spectrum-matching algorithm [47]. The target response spectrum is defined using the Next Generation Attenuation (NGA)-West2 ground motion model [48] and the empirical spectral correlation structure [49]. Each synthetic accelerogram is constructed using a modulated Fourier series with iterative amplitude adjustment to match the target spectrum, followed by baseline correction to remove long-period drift. A total of 1000 synthetic ground motions are generated with a time step of $\Delta t=0.01$ s and a duration of 30 s, resulting in 3,001 time steps per sample. Figure 9 illustrates the finite element model, the response spectra of the synthetic motions, and representative structural responses.
@@ 5.2.2 Prediction performance and discussion
The training, validation, and testing datasets consist of 600, 100, and 300 synthetic ground motion samples, respectively. The neural operator architecture used in this example comprises 8 Fourier layers, 40 retained Fourier modes, and 128 feature channels. The lifting and projection layers follow the configuration used in Section 5.1.2.
@@
Table 3 summarizes the RMSE and relative $\mathcal{L}^{2}$ error across all ten piers, averaged over 10 independent training runs. C-PhysFNO and its linear regression-based refinement consistently outperform the standard FNO baseline. The refinement does not improve the C-PhysFNO prediction, indicating that its primary contribution in this example lies in a preliminary uncertainty quantification for the prediction.
@@ 5.3.1 Structural model and ground motion dataset
This example evaluates the proposed framework using a nine-story steel moment-resisting frame from the SAC Joint Venture benchmark study [50]. The structure is modeled in OpenSees [46], with a total height of 37.19 m and a plan width of 45.73 m. Splice locations for column joints are placed at the first, third, fifth, and seventh stories, each elevated 1.83 m above the beam centerline to accommodate flexural and uplift demands. The foundation is assumed fixed, supported by concrete walls and surrounding soil. Structural responses of interest are the displacement time histories of the nine floors, denoted by $\boldsymbol{u}(t)=\{u_{1}(t),\dots,u_{9}(t)\}$. Key structural properties are summarized in Table 4, and a structural model is shown in Figure 12. The fundamental period of the system is $T=2.27$ s.
The excitation dataset consists of 1,000 recorded ground motions selected from the PEER NGA-West2 database [51]. Selection criteria include: magnitude $7.05.0<M<$ 7.0, Joyner–Boore distance $5015<R_{\text{JB}}<$ 50 km, average shear wave velocity $1800180<V_{S30}<$ 1800 m/s, and uniform time step $\Delta t=0.005$ s. All motions are truncated to a 20-second duration, resulting in 4,001 time steps per sample.
@@ 5.3.2 Prediction performance and discussion
Both the C-PhysFNO and baseline FNO models are trained using 600 samples, with 100 held out for validation and 300 used for testing. Both models employ 8 Fourier layers, 40 retained spectral modes, and 128 feature dimensions. Lifting and projection layers follow the structure described in Section 5.1.2.
Figure 13 illustrates the predicted story-level displacement trajectories produced by C-PhysFNO for three distinct test ground motions. In each panel, the input motion and the corresponding predicted response across all stories are shown. The model demonstrates high accuracy in capturing temporal patterns of nonlinear structural responses. Corresponding refined predictions for top story displacements are shown in the Figure 14, which offers smoother trajectories and quantify uncertainty through posterior intervals. Table 5 summarizes the RMSE and relative $\mathcal{L}^{2}$ errors across all nine stories, averaged over 10 independent training runs. Across all degrees of freedom, C-PhysFNO consistently achieves lower error than the baseline FNO, confirming the benefit of incorporating physics-based intermediate representations for trajectory-level seismic response prediction under recorded ground motion inputs.
@@ 6 Conclusions
Several promising directions remain for future research. First, the integration of adaptive sampling and active learning strategies [52, 53] into the C-PhysFNO framework could further improve data efficiency by prioritizing regions of high predictive uncertainty or critical peak response metrics, especially for probabilistic risk assessment applications. Second, while the present framework assumes deterministic structural parameters, extending it to incorporate epistemic uncertainties in material properties, geometric configurations, and boundary conditions would enable broader applications in structural reliability and design optimization [54, 55]. Finally, although the current framework leverages postprocessing via linear regression to quantify predictive uncertainty, directly integrating uncertainty modeling into the neural operator, such as through Bayesian neural operators [56, 44] could further advance trajectory-level uncertainty quantification.
@@ Acknowledgments
This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00407901).

### v2

@@ 5.1.3 Parametric study
A parametric study is conducted to evaluate neural operator model sensitivity to training dataset size and the number of Fourier kernel layers. Figures 8 and 9 show relative $\mathcal{L}^{2}$ errors across 10 runs for varying configurations. These experiments are conducted for C-PhysFNO using ELS as $\mathcal{H}_{p}$, without postprocessing. While the performance of the standard FNO improves with increased data availability and deeper architectures, the C-PhysFNO consistently achieves lower error across all cases. In fact, the standard FNO achieves competitive accuracy when sufficient training data or model depth is provided, especially in smoother or more structured input–output settings. Notably, the performance gap is most pronounced under data-limited conditions, highlighting the advantage of integrating domain knowledge through physics-based intermediate representations.
@@ 5.2.1 Structural model and ground motion dataset
This example investigates the seismic response of the Auburn Ravine bridge, a prestressed concrete box-girder structure comprising six spans and five bents, each supported by two circular piers konakli2011stochastic . A detailed finite element model is developed in OpenSees mckenna2011opensees , incorporating geometric nonlinearity and material nonlinearities in the pier columns and connection regions. The total bridge length is 166.4 m, and key structural properties are summarized in Table 2. The piers are modeled using force-based beam-column elements with fiber-section discretization. Confined and unconfined concrete fibers and reinforcing steel bars represent axial–bending interaction, while shear and torsion are modeled using uniaxial components. The deck is rigidly connected to the piers, which are fixed at their bases. Translational springs are used to model abutment stiffness. The fundamental period of the bridge is $T=0.38$ s. The quantity of interest is the displacement time history at the top of each of the ten piers, denoted by $\boldsymbol{u}(t)=\{u_{1}(t),\dots,u_{10}(t)\}$.
The ground motion dataset consists of spectrum-compatible synthetic accelerograms generated using the frequency-domain spectrum-matching algorithm yanni2024probabilistic . The target response spectrum is defined using the Next Generation Attenuation (NGA)-West2 ground motion model boore2014nga and the empirical spectral correlation structure baker2008correlation . Each synthetic accelerogram is constructed using a modulated Fourier series with iterative amplitude adjustment to match the target spectrum, followed by baseline correction to remove long-period drift. A total of 1000 synthetic ground motions are generated with a time step of $\Delta t=0.01$ s and a duration of 30 s, resulting in 3,001 time steps per sample. Figure 10 illustrates the finite element model, the response spectra of the synthetic motions, and representative structural responses.
@@ 5.2.2 Prediction performance and discussion
The training, validation, and testing datasets consist of 600, 100, and 300 synthetic ground motion samples, respectively, with model training completed in approximately 40 minutes. The neural operator architecture used in this example comprises 8 Fourier layers, 40 retained Fourier modes, and 128 feature channels. The lifting and projection layers follow the configuration used in Section 5.1.2.
@@
Table 3 summarizes the RMSE and relative $\mathcal{L}^{2}$ error across all ten piers, averaged over 10 independent training runs. C-PhysFNO and its linear regression-based refinement consistently outperform the standard FNO baseline. However, it is worth noting that the standard FNO also achieves reasonably accurate predictions in this example, due in part to the spectral compatibility and smoother structure of the input–output mappings. The refinement does not improve the C-PhysFNO prediction, indicating that its primary contribution in this example lies in a preliminary uncertainty quantification for the prediction.
In addition, Figure 13 demonstrates that both C-PhysFNO and its refinement produce accurate estimates of peak displacements, effectively capturing not only the central tendency but also the overall distribution of the peak response. This performance is likely supported by the use of spectrally compatible input motions, which mitigate frequency mismatch and thus alleviate challenges in the operator learning.
@@
Figure 13: Top-story peak displacement prediction for the Auburn Ravine bridge: (a) scatter plot of predicted versus ground truth peak values across the test set, and (b) estimated probability distribution of peak responses. Both C-PhysFNO and its refinement provide accurate predictions with $R^{2}$ values of 0.9730 and 0.9721, respectively.
@@ 5.3.1 Structural model and ground motion dataset
This example evaluates the proposed framework using a nine-story steel moment-resisting frame from the SAC Joint Venture benchmark study ohtori2004benchmark . The structure is modeled in OpenSees mckenna2011opensees , with a total height of 37.19 m and a plan width of 45.73 m. Splice locations for column joints are placed at the first, third, fifth, and seventh stories, each elevated 1.83 m above the beam centerline to accommodate flexural and uplift demands. The foundation is assumed fixed, supported by concrete walls and surrounding soil. Structural responses of interest are the displacement time histories of the nine floors, denoted by $\boldsymbol{u}(t)=\{u_{1}(t),\dots,u_{9}(t)\}$. Key structural properties are summarized in Table 4, and a structural model is shown in Figure 14. The fundamental period of the system is $T=2.27$ s.
The excitation dataset consists of 1,000 recorded ground motions selected from the PEER NGA-West2 database ancheta2014nga . Selection criteria include: magnitude $5.0<M<7.0$, Joyner–Boore distance $15<R_{\text{JB}}<50$ km, average shear wave velocity $180<V_{S30}<1800$ m/s, and uniform time step $\Delta t=0.005$ s. All motions are truncated to a 20-second duration, resulting in 4,001 time steps per sample.
@@ 5.3.2 Prediction performance and discussion
Both the C-PhysFNO and baseline FNO models are trained using 600 samples, with 100 held out for validation and 300 used for testing, including a total training time of around 60 minutes. Both models employ 8 Fourier layers, 40 retained spectral modes, and 128 feature dimensions. Lifting and projection layers follow the structure described in Section 5.1.2.
Figure 15 illustrates the predicted story-level displacement trajectories produced by C-PhysFNO for three distinct test ground motions. In each panel, the input motion and the corresponding predicted response across all stories are shown. The model demonstrates high accuracy in capturing temporal patterns of nonlinear structural responses. Corresponding refined predictions for top story displacements are shown in the Figure 16, which offers smoother trajectories and quantify uncertainty through posterior intervals. Table 5 summarizes the RMSE and relative $\mathcal{L}^{2}$ errors across all nine stories, averaged over 10 independent training runs. Across all degrees of freedom, C-PhysFNO consistently achieves lower error than the baseline FNO, confirming the benefit of incorporating physics-based intermediate representations for trajectory-level seismic response prediction under recorded ground motion inputs. In addition, Figure 17 further shows that both C-PhysFNO and its refinement accurately capture peak displacement values, aligning closely with the ground truth across the test set.
From a computational perspective, the total offline cost of building the C-PhysFNO model includes both the data generation and model training phases. In this example, generating 600 full nonlinear simulations required approximately 850 minutes on an Intel i9-13900H CPU, while training the neural operator took an additional 60 minutes on the GPU. By contrast, evaluating the same 300 testing samples using full nonlinear analysis would have incurred over 430 minutes of computation time, whereas the trained C-PhysFNO model completes inference for all test cases in under 5 seconds. This highlights a favorable trade-off: while the initial offline cost is nontrivial, the proposed approach enables significant acceleration in downstream evaluations—particularly advantageous for tasks requiring repeated simulations such as uncertainty quantification, design optimization, or real-time decision support. Such computational efficiency positions the proposed method as a viable surrogate for forward modeling in many-query settings. However, for limited-scope analyses involving only a small number of forward simulations, conventional numerical solvers may remain preferable. These practical considerations reinforce the intended use case of the proposed framework and complement earlier discussions on predictive accuracy and operator generalization.
@@
Figure 17: Top-story peak displacement prediction for the nine-story steel moment-resisting frame: (a) scatter plot of predicted versus ground truth peak values across the test set, and (b) estimated probability distribution of peak responses. Both C-PhysFNO and its refinement provide accurate predictions with $R^{2}$ values of 0.9920 and 0.9914, respectively.
@@ 6 Conclusions
Several promising directions remain for future research. First, the integration of adaptive sampling and active learning strategies kim2024adaptive , zhou2025multi into the C-PhysFNO framework could further improve data efficiency by prioritizing regions of high predictive uncertainty or critical peak response metrics, especially for probabilistic risk assessment applications. Second, while the present framework assumes deterministic structural parameters, extending it to incorporate epistemic uncertainties in material properties, geometric configurations, and boundary conditions would enable broader applications in structural reliability and design optimization celik2010seismic , kim2024active . Finally, although the current framework leverages postprocessing via linear regression to quantify predictive uncertainty, directly integrating uncertainty modeling into the neural operator, such as through Bayesian neural operators magnani2022approximate , bulte2025probabilistic could further advance trajectory-level uncertainty quantification.
@@ Acknowledgments
This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00407901). This work was also supported by the faculty research fund of Sejong University in 2025. The authors also thank the two anonymous reviewers for their thoughtful and constructive comments, which greatly improved the quality of this work.